A boy rows his boat 8 km downstream in 1 hr and 4 mins and return in 2 and 2/7hr.At what rate he will row in still water.What is the rate of the current.
Answer by checkley77(12844)

8=1.06667r
r=8/1.06667
r=7.5 kph.
8=2.2857r
r=8/2.2857
r=3.5 kph.
(7.5-3.5)/2=4/2=2 kph. for the current.
7.5-2=5.5 kph. for the rowing speed.
3.5+2=5.5 ditto.